OpenStudy (anonymous):

2. What is the value of the expression 1/3^(-5) A. 243 B.15 C.1/15 D.1/243 3. What is the value of the expression 3/ 3^(-2) A.27 B.18 C.1/2 D.1/3 4. What value of K ^(-3) = 1/27 A.-81 B.-9 C.-3 D.3 5. What value of n solves the expression 3^(n) = 1/81 A.-243 B.-27 C.-4 D.4 Please help

OpenStudy (johnweldon1993):

\[\large \frac{1}{3^{-5}} = 3^5 = ?\]

OpenStudy (johnweldon1993):

\[\large \frac{3}{3^{-2}} = 3\times 3^2 = ?\]

OpenStudy (johnweldon1993):

So basically whenever you have a negative exponent, you flip it to the other side of the fraction and turn it positive